OpenStudy (anonymous):

please distribute and explain 3(x+1)+2(x-4) WILL MEDAL AND FAN!!!!!

OpenStudy (brrandyn):

Stargurl, multiply coefficient outside the parenthesis. 3 * x + 3*1 + 2x -8

OpenStudy (welshfella):

first distribute the 3 over the x + 1 in the first parentheses This means you multiply the x than the + 2 by 3:- 3(x + 1) = 3*x + 3*1 = 3x + 3

OpenStudy (brrandyn):

Then add :) 3x + 3 + 2x -8 5x -5 is your solution.

OpenStudy (welshfella):

- then do the same process with + 2(x - 4) (remember that a -quantity times a + = - quantity)

OpenStudy (anonymous):

@welshfella still dont understand it like every time i do this i know but then i forget an it gets harder every time someone tries to explain it again

OpenStudy (anonymous):

ohhh i get it thank you @Brrandyn

OpenStudy (welshfella):

the best way to get confident is to practice them